# Some examples of use of transfinite induction in analysis

### v1

@@ 1 Introduction
In this paper we want to advertise a different kind of argument, that somehow replaces these ‘big steps’ with possibly ‘small ones’. It relies on the fact that any monotone map from the first uncountable ordinal to the reals must be eventually constant, thus if we index our iterative procedure over the ordinals and at every step we try to augment our given real valued function, then regardless of how big our steps are, we are sure that after a countable number of steps we are done anyway.
@@
With all this being said, we shall also provide an alternative proof of the existence of a Maximal Globally Hyperbolic Development that has nothing to do with transfinite recursion and ordinals. Rather, we construct a way of quantifying the size of a development, so that a maximal one can be built via ‘big steps’, along classical lines. Still, the fact that this quantification is not-so-evidently found is perhaps an argument in favour of the relevance of the line of reasoning discussed here in the mathematical conversation.
I thought for the first time at a ‘small steps’ argument via transfinite recursion of the kind just discussed in my PhD thesis [4, Proposition 5.10]. I mentioned this to a number of colleagues and to all of them the argument seemed new; still, I cannot exclude that in circles far from mine this is a well-known line of reasoning. In [4] I needed a statement closely resembling Ekeland’s variational principle, of which I was unaware, and admittedly I did not find a way to make a proper quantification, thus I resorted on this other approach. Later, any time I was tempted to use this line of thought, a quantification was easily found, making the small-steps argument irrelevant. I thus never published anything on the matter, limiting myself to asking questions on the matter on MathOverflow [1], especially in connection to the kind of Axiom of Choice needed to carry out the argument.
@@ 2 Abstract presentation of the mechanism
The next lemma gives an abstract presentation of the argument we are going to use:
Lemma 2.1 (Transfinite recursion over countable ordinals).
Let $A$ be a set, $F\subset A$ a subset and $\mathcal{S}$ a non-empty collection of sequences of length $\omega_{1}$ with the following properties:
@@
    There is no $(a_{\alpha})_{\alpha<\omega_{1}}$ such that for every $\alpha<\omega$ we have $(a_{\alpha^{\prime}})_{\alpha^{\prime}<\alpha}\in\mathcal{S}$.
@@
Define recursively a transfinite sequence as follows. Let $a_{0}\in A$ be arbitrary. Then let $\alpha<\omega_{1}$ and assume to have already defined $a_{\alpha^{\prime}}$ for every $\alpha^{\prime}<\alpha$} and that $(a_{\alpha^{\prime\prime}})_{\alpha^{\prime\prime}\leq\alpha^{\prime}}\in\mathcal{S}$ for every $\alpha^{\prime}<\alpha$. Then $(a_{\alpha^{\prime}})_{\alpha^{\prime}<\alpha}\in\mathcal{S}$: this is obvious if $\alpha$ is a successor, while if it is limit we use $(ii)$. If for some $\alpha^{\prime}<\alpha$ we have $a_{\alpha^{\prime}}\in F$ we are done, otherwise by $(i)$ there is $a_{\alpha}\in A$ such that $(a_{\alpha^{\prime}})_{\alpha^{\prime}\leq\alpha}\in\mathcal{S}$.
@@
We might replace assumption $(iii)$ above with the assumption that sequences in $\mathcal{S}$ are injective, as this is what happens in the applications we have in mind. However, for us it might be relevant to know that the recursive procedure only involves countable ordinals (e.g. in discussing the application to the Hahn-Jordan decomposition we need to be sure that the sets we produce at each step are measurable) and for this assumption $(iii)$ matters. A further advantage of having it is that it clarifies which version of Axiom of Choice is needed to carry on the argument: as Prof. A. Karagila patiently explained me, Lemma 2.1 is equivalent to ${\rm DC}_{\omega_{1}}$.
@@ 3.3 Maximal Globally Hyperbolic Development
    $(M,g)$ is a Ricci-flat spacetime (and thus solves Einstein’s equations in the vacuum),
@@
In particular, $M$ admits a Cauchy hypersurface and thus is globally hyperbolic. For the terminology we refer to [2]. We observe that the embedding $\varphi:\Sigma\to M$ induces, for every $p\in\Sigma$, a unique causal isometry $\Phi_{p}$ from $(\mathbb{R}\times T_{p}\Sigma,{\rm d}t^{2}-h)$ to $(T_{\varphi(p)}M,g)$ sending $(0,v)$ to ${\rm d}\varphi(v)$, being intended that $(1,0)\in\mathbb{R}\times T_{p}\Sigma$ is future directed.
@@
Proof via ‘small steps’ The uniqueness claim is a direct consequence of the definition of maximality and the fact that $\preceq$ is a partial order on (isomrphism classes of) developments: if $M,\tilde{M}$ are developments satisfying $(1)$, then $\tilde{M}\preceq M$ and $M\preceq\tilde{M}$, proving that $\tilde{M}\cong M$.
The existence proof will be split in two parts. In the first part we show existence of a development $(M,g,\varphi)$ that is maximal in the sense, weaker than $(1)$, that:
  • 1’)
    (Maximality) If $(M^{\prime},g^{\prime},\varphi^{\prime})$ is another development and $M\preceq M^{\prime}$ then $M^{\prime}\cong M$.
In the second part we show that any such $M$ is also maximal in the sense of $(1)$.
We start with the first part. Let $A$ be the collection of all developments of $(\Sigma,h,\kappa)$. We avoid the usual set-theoretic issues related to the “all” above by asking the underlying set on which the manifold structures are given to be a fixed set of cardinality of continuum (notice that Proposition 3.4 ensures that this is the cardinality of any connected manifold equipped with a smooth metric tensor). Let $F\subset A$ be the collection of all the developments satisfying $(1^{\prime})$ above. Let also $\mathcal{S}$ be the collection of $\alpha$-sequences $(M_{\beta})_{\beta<\alpha}$ in $A$, $\alpha<\omega_{1}$, that are strictly increasing, meaning that for any $\beta_{1}<\beta_{2}<\alpha$ we have $M_{\beta_{1}}\preceq M_{\beta_{2}}$ and $M_{\beta_{2}}\not\preceq M_{\beta_{1}}$. We show that $A,F,\mathcal{S}$ satisfy the assumptions in Lemma 2.1.
@@
It follows by Lemma 2.1 that $F$ is not empty, i.e. that there is a development satisfying $(1^{\prime})$, as desired.
We turn to the second part of the proof. We want to show that a development as in $(1^{\prime})$ is maximal also in the sense of $(1)$ in the statement. Inspecting the proof in [2], we see that the axiom of Choice is used, via Zorn’s lemma, to prove existence of a ‘maximal common development’ in the following sense:
[…]
In our line of thought, the existence of this maximal common development follows along the very same arguments just used: we let $A$ be the collection of developments, $F$ that of those satisfying the maximality property just stated and $\mathcal{S}$ the collection of strictly increasing $\alpha$-sequences in $A$ bounded from above by both $M_{1}$ and $M_{2}$. The non-emptiness of $\mathcal{S}$ follows by Theorem 3.3 and then that of $F$ by Lemma 2.1.
Once ‘maximal common development’ $M$ in the sense above is built, the gluing argument given in [2] shows that it is $\preceq$-maximal among all the common developments of $M_{1}$ and $M_{2}$ (not just among those comparable with $M$) and then the proof can continue as in [2]. $\square$
@@
and let $(t_{n},v_{n})_{n\in\mathbb{N}}$ be an enumeration of $D^{T}$. Also, let us fix an order siomorphism $\eta:[0,+\infty]\to[0,1]$.
Let $(M,g,\varphi)$ be a development of $(\Sigma,h,\kappa)$, recall that $\varphi$ induces, for any $p\in\Sigma$, an isomorphism $\Phi$ of $(\mathbb{R}\times T_{p}\Sigma,dt^{2}-h)$ and $(T_{\varphi(p)}M,g)$ sending $(1,0)$ to a future vector. For each $n\in\mathbb{N}$, $\Phi(t_{n},v_{n})$ is a future timelike vector in $M$: let $\gamma_{n}=\gamma_{n}(M)$ be the maximal geodesics (intended as solution of the geodesic equation) in $M$ with $\gamma_{n}^{\prime}(0)=\Psi(t_{n},v_{n})$. Then smoothness of $M$ grants that this is a good definition. Write $(-a_{n},b_{n})\subset\mathbb{R}$ with $a_{n},b_{n}>0$ for the interval of definition of $\gamma_{n}$. Then define
$F(M)=F(M,g,\varphi):=\sum_{n\in\mathbb{N}}\tfrac{1}{2^{n}}(\eta(b_{n})+\eta(a_{n}))\ \in[0,4].$
By definition it is obvious that $F$ is monotone,, i.e. that if $(M_{1},g_{1},\varphi_{1})$ and $(M_{2},g_{2},\varphi_{2})$ are two developments with $M_{1}\preceq M_{2}$ then $F(M_{1})\leq F(M_{2})$ (because the domain of $\gamma_{n}(M_{1})$ is contained in that of $\gamma_{n}(M_{2})$ for every $n\in\mathbb{N}$). We now observe that $F$ is strictly monotone, i.e.
@@
Having established the strict monotonicity (3.6) we proceed as follows. The uniqueness property $(2)$ is obvious from the fact that $\preceq$ is a partial order. For the maximality $(1)$ we first prove the weaker claim
  • 1’)
    (Maximality) If $(M^{\prime},g^{\prime},\varphi^{\prime})$ is another development and $M\preceq M^{\prime}$ then $M^{\prime}\cong M$.
Given a development $(M,g\varphi)$ we define $G(M)\in[0,4]$ as
@@
and notice that $M_{1}\preceq M_{2}$ implies $G(M_{1})\geq G(M_{2})$. Now define recursively a sequence $(M_{n})$ of developments as follows. Let $(M_{0},g_{0},\varphi_{0})$ be an arbitrary development as given by Theorem 3.3. Having defined $(M_{n},g_{n},\varphi_{n})$ pick $(M_{n+1},g_{n+1},\varphi_{n+1})$ so that $M_{n}\preceq M_{n+1}$ and $F(M_{n+1})\geq\tfrac{1}{2}(F(M_{n})+G(M_{n}))$. The definition of $G(M)$ ensures that such $M_{n+1}$ exists. Gluing all these developments (as in [2]), we find a development $(M,g,\varphi)$ with $M_{n}\preceq M$ for every $n\in\mathbb{N}$. The construction ensures that $G(M_{n+1})-F(M_{n+1})\leq G(M_{n})-\tfrac{1}{2}(F(M_{n})+G(M_{n}))\leq G(M_{n})-F(M_{n})$ and thus $G(M_{n})-F(M_{n})\leq\tfrac{1}{2^{n}}(G(M_{0})-F(M_{0}))\to 0$, so that the monotonicities of $F,G$ imply that $G(M)=F(M)$, which by (3.6) means that $M$ is maximal as in $(1^{\prime})$.
To conclude the proof we now need to show that a development maximal as in $(1^{\prime})$ is so as in $(1)$. Arguing as in the previous proof (actually: as in [2]) this will follow if we prove the existence of a ‘maximal common development’ of two given ones. This can be achieved along the very same lines just used, by quantifying the size of common developments using $F$. We omit the details. $\square$

### v2

@@ 1 Introduction
In this paper we want to advertise a different kind of argument, that somehow replaces these ‘big steps’ with possibly ‘small ones’. It relies on the fact that any monotone map from the first uncountable ordinal to the reals must be eventually constant (we recall the proof of this basic fact in Proposition A.1), thus if we index our iterative procedure over the ordinals and at every step we try to augment our given real valued function, then regardless of how big our steps are, we are sure that after a countable number of steps we are done anyway.
@@
With all this being said, we shall also provide an alternative proof of the existence of a Maximal Globally Hyperbolic Development that has nothing to do with transfinite recursion and ordinals. Rather, we construct a way of quantifying the size of a development, see (3.7), so that a maximal one can be built via ‘big steps’, along classical lines. Still, the fact that this quantification is not-so-evidently found is perhaps an argument in favour of the relevance of the line of reasoning discussed here in the mathematical conversation.
In conclusion, the main takeaways of this paper are:
  1. 1.
    The illustration of the potential use of transfinite induction/recursion over countable ordinals to deduce ‘honest’ results in geometric analysis. The advantage, if any, over more standard arguments is that it permits to pay less attention to the choices made during the induction procedure. We stress that this line of thought is based over ${\sf DC}_{\omega_{1}}$ (i.e. Dependent Choice indexed over the first uncountable ordinal $\omega_{1}$) that in turn is still a Choice strong enough to imply, for instance, existence of non-measurable subsets of the reals.
  2. 2.
    A modification of the original proof of existence of Maximal Global Hyperbolic Developments, based on the quantification (3.7). This has nothing to do with ordinals and transfinite induction and because of this might be of some interest for the mathematical GR community.
    The use of the quantification (3.7) eliminates the need of Zorn’s lemma replacing it with the more commonly used Countable Dependent Choice. In this sense this part of the work provides an alternative way to ‘dezornify’ the original proof in [3], comparable to that given in [9]. We remark that in [11] a further different proof has been given, that relies on no Choice at all for what concerns the construction part of MGHD. Our proof, much like several other proofs by iteration, does not allow to fully drop Choice and we shall often use the usual Countable Dependent Choice throughout the text. We do not regard this as a weakness of our approach, as we believe anyway – in line with [2] – that CDC is crucially needed in most of the aspects of (geometric) analysis.
I thought for the first time at a ‘small steps’ argument via transfinite recursion of the kind just discussed in my PhD thesis [5, Proposition 5.10]. I mentioned this to a number of colleagues and to all of them the argument seemed new; still, I cannot exclude that in circles far from mine this is a well-known line of reasoning. In [5] I needed a statement closely resembling Ekeland’s variational principle, of which I was unaware, and admittedly I did not find a way to make a proper quantification, thus I resorted on this other approach. Later, any time I was tempted to use this line of thought, a quantification was easily found, making the small-steps argument irrelevant. I thus never published anything on the matter, limiting myself to asking questions on the matter on MathOverflow [1], especially in connection to the kind of Axiom of Choice needed to carry out the argument.
@@
I also want to thank the referees for the careful reading of the manuscript and their useful suggestions and also Prof. A. Lerario for having brought to my attention the Prüfer surface, relevant in Remark 3.7.
@@ 2 Abstract presentation of the mechanism
Here we present the abstract mechanism behind our use of transfinite induction. We assume the reader familiar with the basic theory of ordinals, as can be found for instance in [6]. An informal, but rigorous, presentation of all the necessary ingredients is also given in the appendix.
[…]
The technique we want to discuss is for proving existence results. So we are given a non-empty set $A$ and a subset $F\subset A$ of ‘final’ elements that we intend to show to be non-empty. The way we have to probing that $F$ is not empty is via a set $\mathcal{S}$ of sequences in $A$ that, roughly said, have the property that ‘if a sequence in $\mathcal{S}$ does not meet $F$, then it can be extended’. If this occurs, then indeed at least one sequence in $\mathcal{S}$ must meet $F$, which therefore is not empty. Even more: if we have an upper bound on the length of sequences in $\mathcal{S}$ (in our case this is $\omega_{1}$), then $F$ must be met before reaching such upper bound.
We emphasize that for this argument to work it is imperative to work with sequences indexed by ordinals, as easy examples show that ordinary sequences cannot lead to the same conclusion.
Lemma 2.1.
Let $A$ be a set, $F\subset A$ a subset and $\mathcal{S}$ a non-empty collection of sequences in $A$ of length $\omega_{1}$ with the following properties:
@@
    There is no $(a_{\alpha})_{\alpha<\omega_{1}}$ such that for every $\alpha<\omega_{1}$ we have $(a_{\alpha^{\prime}})_{\alpha^{\prime}<\alpha}\in\mathcal{S}$.
@@
Define recursively a transfinite sequence as follows. Let $(a_{\alpha^{\prime}})_{\alpha^{\prime}<\alpha_{0}}\in\mathcal{S}$ be arbitrary. If this satisfies the conclusion we are done. Otherwise we look for extensions in $\mathcal{S}$ of this sequence. Then let $\alpha<\omega_{1}$ be with $\alpha_{0}<\alpha$ and assume to have already defined $a_{\alpha^{\prime}}$ for every $\alpha_{0}\leq{\alpha^{\prime}<\alpha}$ and that $(a_{\alpha^{\prime\prime}})_{\alpha^{\prime\prime}\leq\alpha^{\prime}}\in\mathcal{S}$ for every $\alpha^{\prime}<\alpha$. Then $(a_{\alpha^{\prime}})_{\alpha^{\prime}<\alpha}\in\mathcal{S}$: this is obvious if $\alpha$ is a successor, while if it is limit we use $(ii)$. If for some $\alpha^{\prime}<\alpha$ we have $a_{\alpha^{\prime}}\in F$ we are done, otherwise by $(i)$ there is $a_{\alpha}\in A$ such that $(a_{\alpha^{\prime}})_{\alpha^{\prime}\leq\alpha}\in\mathcal{S}$.
@@
We might replace assumption $(iii)$ above with the assumption that sequences in $\mathcal{S}$ are injective, as this is what happens in the applications we have in mind. However, for us it might be relevant to know that the recursive procedure only involves countable ordinals (e.g. in discussing the application to the Hahn-Jordan decomposition we need to be sure that the sets we produce at each step are measurable) and for this assumption $(iii)$ matters. A further advantage of having it is that it clarifies which version of Axiom of Choice is needed to carry on the argument: as Prof. A. Karagila patiently explained me, Lemma 2.1 is equivalent to ${\rm DC}_{\omega_{1}}$, i.e. Dependent Choice indexed over $\omega_{1}$ (we refer to the comprehensive monograph [7], and in particular Chapter 8 in there, for more on the matter. We illustrate this by first recalling what ${\rm DC}_{\omega_{1}}$ says. Recall that given a set $A$, by $A^{<\omega_{1}}$ it is meant the collection of $\alpha$-sequences in $A$ for $\alpha<\omega_{1}$.
Axiom (${\sf DC}_{\omega_{1}}$).
Let $A$ be a non-empty set and $G$ a map from $A^{<\omega_{1}}$ to non-empty subsets of $A$. Then there exists an $\omega_{1}$-sequence $f:\omega_{1}\to A$ such that
[…]
holds for every $\alpha<\omega_{1}$.
[…]
Proposition 2.2.
In Zermelo-Fraekel set theory, Lemma 2.1 is equivalent to Axiom ${\sf DC}_{\omega_{1}}$.
Proof.
Lemma 2.1 implies Axiom ${\sf DC}_{\omega_{1}}$. We argue by contradiction. We thus assume that ${\sf DC}_{\omega_{1}}$ does not hold, i.e. that for $A$ and $G$ as in the statement of the axiom there is no $f:\omega_{1}\to A$ such that (2.1) holds for every $\alpha<\omega_{1}$.
Let $F$ be the empty set and $\mathcal{S}$ be the collection of sequences $f$ valued in $A$ such that (2.1) holds for every $\alpha$ smaller than the length of the sequence. We verify that $A,F,\mathcal{S}$ verify $(i),(ii),(iii)$ in Lemma 2.1. Property $(i)$ is obvious from the fact that $G$ takes values in non-empty subsets of $A$, $(ii)$ is obvious by definition and $(iii)$ is our assumption of failure of ${\sf DC}_{\omega_{1}}$. We thus see that Lemma 2.1 fails, because the data satisfy the assumption but the set $F$ is empty, contradicting the conclusion.
Axiom ${\sf DC}_{\omega_{1}}$ implies Lemma 2.1. Let $A,F,\mathcal{S}$ be as in the statement of Lemma 2.1. Let $\mathcal{S}^{\prime}$ be the collection of sequences in $\mathcal{S}$ and of their restriction to ordinals smaller than their original lengths. Let us pick one extra element, call it $\bar{a}$, not belonging to $A$ and let us define $\bar{A}:=A\cup\{\bar{a}\}$ and $\bar{F}:=F\cup\{\bar{a}\}$. We shall think $\mathcal{S},\mathcal{S}^{\prime}$ as set of maps from countable ordinals into $\bar{A}$. Now for $f\in\bar{A}^{<\omega_{1}}$ we define $G(f)$ as follows: if $f$ is in $\mathcal{S}^{\prime}$ and the image does not meet $F$, then $G(f)\subset A\subset\bar{A}$ is the collection of those values so that if we extend $f$ one further step via this value, the extension is still in $\mathcal{S}^{\prime}$ (notice that by definition of $\mathcal{S}^{\prime}$ and assumption $(i)$ in Lemma 2.1, in this case $G(f)$ is not empty). In all the other cases we put $G(f):=\{\bar{a}\}$.
[…]
@@ 3.3 Maximal Globally Hyperbolic Development
    $(M,g)$ is a Ricci-flat time-oriented spacetime,
@@
In particular, $M$ solves Einstein’s equations in the vacuum. Also, it admits a Cauchy hypersurface and thus is globally hyperbolic. For the terminology we refer to [3]. We observe that the embedding $\varphi:\Sigma\to M$ induces, for every $p\in\Sigma$, a unique causal isometry $\Phi_{p}$ from $(\mathbb{R}\times T_{p}\Sigma,{\rm d}t^{2}-h)$ to $(T_{\varphi(p)}M,g)$ sending $(0,v)$ to ${\rm d}\varphi(v)$, being intended that $(1,0)\in\mathbb{R}\times T_{p}\Sigma$ is future directed and is therefore sent to the future-directed normal to $S$ at $\varphi(p)$.
@@
Preliminary considerations The uniqueness claim is a direct consequence of the definition of maximality and the fact that $\preceq$ is a partial order on (isomrphism classes of) developments: if $M,\tilde{M}$ are developments satisfying $(1)$, then $\tilde{M}\preceq M$ and $M\preceq\tilde{M}$, proving that $\tilde{M}\cong M$.
For existence, let us briefly recall how the original proof in [3] works:
  • A)
    One proves that there exists a development $(M,g,\varphi)$ that is maximal in the sense, weaker than $(1)$, that:
    if $(M^{\prime},g^{\prime},\varphi^{\prime})$ is another development and $M\preceq M^{\prime}$ then $M^{\prime}\cong M$. (3.6)
    Here the authors of [3] use Zorn’s lemma: it is easy to see that any chain of developments admits a maximizer (just ‘glue’ all the developments together to create one bigger than any of those in the chain), hence Zorn’s lemma ensures the existence of the desired maximizer.
  • B)
    One proves that given two developments $M_{1},M_{2}$ there exists a development $M$ that is $\preceq M_{1}$ and $\preceq M_{2}$ and is maximal in the sense that:
    if $N$ is another development $\preceq M_{1}$, $\preceq M_{2}$ and $M\preceq N$, then $N\cong M$.
    Even in this case, in [3] the argument makes use of Zorn’s lemma by observing that any chain of ‘common developments’ admits a maximizer that is still a common development.
  • C)
    One proves that if $M_{1}$ and $M_{2}$ are two developments and $M$ is a maximal common development in the sense of $\sf(B)$} above, then gluing $M_{1}$ and $M_{2}$ along $M$ produces a development, that obviously is $\succeq M_{1}$ and $\succeq M_{2}$. Here the difficult part, is in showing that the result of the gluing is Hausdorff. Very shortly and roughly said, if it were not there would $p_{1}\in M_{1}$ and $p_{2}\in M_{2}$ that do no possess disjoint neighbourhoods in the gluing. It is easy to see that this can occur only if they belong to respective boundaries of the glued regions, but then - with some work that relies on local uniqueness of solutions of Einstein’s equations - one can show that the common development can be extended to incorporate also $p_{1},p_{2}$, contradicting maximality.
  • D)
    The conclusion is now easy: we know that there is a development $M$ that is maximal in the sense of $\sf(A)$}. Let $M^{\prime}$ be another development, $N$ a maximal common development as in $\sf(B)$} and use $\sf(C)$} to glue $M$ and $M^{\prime}$ along $N$. This new development is $\succeq M$ and thus, by maximality of $M$, must be isomorphic to $M$. It is easy to see that for this to happen we must have that $N$ is the whole $M^{\prime}$, i.e. that $M^{\prime}\preceq M$. In other words, $M$ is the desired development maximal in the sense $(1)$ of the statement.
Points $\sf(C),(D)$} in this argument do not require any Choice at all, so we won’t give further details and refer rather to the original source and to the comprehensive text [8]. We shall instead focus on how the ‘small steps’ and ‘big steps’ arguments can be used to tackle points $\sf(A),(B)$}.
Proof via ‘small steps’ We start with proving point $\sf(A)$}. Let $A$ be the collection of all developments of $(\Sigma,h,\kappa)$. We avoid the usual set-theoretic issues related to the “all” above by asking the underlying set on which the manifold structures are given to be a fixed set of cardinality of continuum (notice that Proposition 3.4 ensures that this is the cardinality of any connected manifold equipped with a smooth metric tensor). Let $F\subset A$ be the collection of all the developments satisfying (3.6) above. Let also $\mathcal{S}$ be the collection of $\alpha$-sequences $(M_{\beta})_{\beta<\alpha}$ in $A$, $\alpha<\omega_{1}$, that are strictly increasing, meaning that for any $\beta_{1}<\beta_{2}<\alpha$ we have $M_{\beta_{1}}\preceq M_{\beta_{2}}$ and $M_{\beta_{2}}\not\preceq M_{\beta_{1}}$. We show that $A,F,\mathcal{S}$ satisfy the assumptions in Lemma 2.1.
@@
It follows by Lemma 2.1 that $F$ is not empty, i.e. that there is a development satisfying (3.6), as desired.
We turn to point $\sf(B)$}. We want to show that a development as in (3.6) is maximal also in the sense of $(1)$ in the statement. This follows along the very same arguments just used: we let $A$ be the collection of developments, $F$ that of those satisfying the maximality property just stated and $\mathcal{S}$ the collection of strictly increasing $\alpha$-sequences in $A$ bounded from above by both $M_{1}$ and $M_{2}$. The non-emptiness of $\mathcal{S}$ follows by Theorem 3.3 and then that of $F$ by Lemma 2.1. $\square$
@@
and let $(t_{n},v_{n})_{n\in\mathbb{N}}$ be an enumeration of $D^{T}$. Also, let us fix an order isomorphism $\eta:[0,+\infty]\to[0,1]$.
Let $(M,g,\varphi)$ be a development of $(\Sigma,h,\kappa)$, recall that $\varphi$ induces, for any $p\in\Sigma$, an isomorphism $\Phi$ of $(\mathbb{R}\times T_{p}\Sigma,dt^{2}-h)$ and $(T_{\varphi(p)}M,g)$ sending $(1,0)$ to a future vector. For each $n\in\mathbb{N}$, $\Phi(t_{n},v_{n})$ is a future timelike vector in $M$: let $\gamma_{n}=\gamma_{n}(M)$ be the maximal geodesics (intended as solution of the geodesic equation) in $M$ with $\gamma_{n}^{\prime}(0)=\Phi(t_{n},v_{n})$. Then smoothness of $M$ grants that this is a good definition. Write $(-a_{n},b_{n})\subset\mathbb{R}$ with $a_{n},b_{n}>0$ for the interval of definition of $\gamma_{n}$. Then define
$F(M)=F(M,g,\varphi):=\sum_{n\in\mathbb{N}}\tfrac{1}{2^{n}}(\eta(b_{n})+\eta(a_{n}))\ \in[0,4].$ (3.7)
By definition it is obvious that $F$ is monotone, i.e. that if $(M_{1},g_{1},\varphi_{1})$ and $(M_{2},g_{2},\varphi_{2})$ are two developments with $M_{1}\preceq M_{2}$ then $F(M_{1})\leq F(M_{2})$ (because the domain of $\gamma_{n}(M_{1})$ is contained in that of $\gamma_{n}(M_{2})$ for every $n\in\mathbb{N}$). We now observe that $F$ is strictly monotone, i.e.
@@
Having established the strict monotonicity (3.8) we proceed by proving points $\sf(A),(B)$}. We start with $\sf(A)$}. Given a development $(M,g,\varphi)$ we define $G(M)\in[0,4]$ as
@@
and notice that $M_{1}\preceq M_{2}$ implies $G(M_{1})\geq G(M_{2})$. Now define recursively a sequence $(M_{n})$ of developments as follows. Let $(M_{0},g_{0},\varphi_{0})$ be an arbitrary development as given by Theorem 3.3. Having defined $(M_{n},g_{n},\varphi_{n})$ pick $(M_{n+1},g_{n+1},\varphi_{n+1})$ so that $M_{n}\preceq M_{n+1}$ and $F(M_{n+1})\geq\tfrac{1}{2}(F(M_{n})+G(M_{n}))$. The definition of $G(M)$ ensures that such $M_{n+1}$ exists. Gluing all these developments (as in [3]), we find a development $(M,g,\varphi)$ with $M_{n}\preceq M$ for every $n\in\mathbb{N}$. The construction ensures that $G(M_{n+1})-F(M_{n+1})\leq G(M_{n})-\tfrac{1}{2}(F(M_{n})+G(M_{n}))\leq\tfrac{1}{2}(G(M_{n})-F(M_{n}))$ and thus $G(M_{n})-F(M_{n})\leq\tfrac{1}{2^{n}}(G(M_{0})-F(M_{0}))\to 0$, so that the monotonicities of $F,G$ imply that $G(M)=F(M)$, which by (3.8) means that $M$ is maximal in the sense of (3.6).
For $\sf(B)$} we can argue along the lines just used by quantifying the size of common developments using $F$. We omit the details. $\square$